# A Survey on Vision-Language-Action Models for Embodied AI

### v4

@@ V-F Embodied Question Answering Benchmarks
Table VII: Embodied question answering benchmarks. Acc: accuracy. PPL: perplexity.
# QA # Videos Video source Answer type Active Data collection Metrics
EQA [223] 5K 750 envs House3D simulator Answer set 172 Yes Template Acc
IQUAD [224] 75K 30 rooms AI2-THOR Multiple-choice Yes Template Acc
MT-EQA [225] 19.3K 588 envs House3D simulator Binary answer Yes Template Acc
MP3D-EQA [226] 1,136 83 envs MatterPort3D Answer set 53 Yes Template Acc
EgoVQA [227] 600 16 IU Multi-view Multiple-choice (1 out of 5) No Human annotators Acc
EgoTaskQA [228] 40K 2K LEMMA Answer set (open-answer and binary verifications) No Human annotators for relationship triples, then template Acc
EQA-MX [229] 8.2M 750K images CAESAR simulator Answer set No Question templates, answer set (simulator ground truth) Acc
OpenEQA [230] 557 + 1079 180 envs HM3D + ScanNet Open-answer (LLM scorer) No Human annotators Score
@@ VI Challenges and Future Directions
Multi-Modality
VLAs inherit many challenges associated with multimodal models, such as obtaining useful embeddings and aligning different modalities. Current approaches, like ImageBind [231] and LanguageBind [232], align different modalities to the image or language embedding space, respectively. Within a unified embedding space, MLLMs can accommodate diverse modalities and become more general. However, whether focusing on embeddings alone is sufficient remains under debate. Although beyond the scope of VLAs, other modalities—such as audio [60] and haptics [233]—have proven useful for certain embodied AI applications. Incorporating additional modalities is often desirable, but they also introduce extra complexity into the model design.
@@
Robotics has always raised various ethical, legal, and societal concerns. These include risks related to privacy, safety, job displacement, bias in decision-making, and the impact on social norms and human relationships.
@@
Most current VLAs focus on household and industrial settings [2], which shape the available datasets and models. Given the potential of VLA models, a wide range of new applications is possible. One particularly important application is in healthcare, such as surgical robots [234] and care robots [235]. Healthcare demands higher safety and privacy standards and may necessitate novel techniques such as human-in-the-loop (HITL) control and federated learning. Moreover, due to the significant domain gap, specialized vision models might be needed for medical images.
@@ VII Conclusion
Vision-language-action models hold immense promise for enabling embodied agents to interact with the physical world and fulfill users’ instructions. This paper is the first survey to review large VLAs alongside other generalized VLAs. Our taxonomy provides a high-level overview of three main lines of research: key components, control policies, and task planners. We meticulously summarize and compare their model designs and technical details. Additionally, we highlight essential resources for training and evaluating VLAs, such as datasets, simulators, and benchmarks. We hope this survey captures the rapidly evolving landscape of embodied AI and inspires future research.

### v5

@@ V-F Embodied Question Answering Benchmarks
Table VII: Embodied question answering benchmarks. Explore: active exploration.
Benchmark QA Pairs Scenes Source Answer Type Collection Explore Metrics
EQA [197] 5K 750 envs House3D simulator Answer set (172 answers) Template Accuracy
IQUAD [198] 75K 30 rooms AI2-THOR Multiple choice Template Accuracy
MT-EQA [199] 19.3K 588 envs House3D simulator Binary answer Template Accuracy
MP3D-EQA [200] 1,136 83 envs MatterPort3D Answer set (53 answers) Template Accuracy
EgoVQA [201] 600 16 videos IU Multi-view Multiple choice (1 out of 5) Human annotators Accuracy
EgoTaskQA [202] 40K 2K videos LEMMA dataset Open answer, binary verification Human annotators & template Accuracy
EQA-MX [203] 8.2M 750K images CAESAR simulator Answer set Question templates, answer set Accuracy
OpenEQA [204] 557 + 1079 180 envs HM3D + ScanNet Open answer Human annotators LLM Score
@@ VI Challenges and Future Directions
Multimodality
VLAs inherit many challenges associated with multimodal models, such as obtaining useful embeddings and aligning different modalities. Current approaches, like ImageBind [205] and LanguageBind [206], align different modalities to the image or language embedding space, respectively. Within a unified embedding space, MLLMs can accommodate diverse modalities and become more general. However, whether focusing on embeddings alone is sufficient remains under debate. Although beyond the scope of VLAs, other modalities—such as audio [47], haptics [207], and gaze [208]—have proven useful for certain embodied AI applications. For instance, modeling human gaze data using an additional gaze network [209] or incorporating it via an auxiliary loss [210] has been shown to enhance the performance of the primary policy network. These approaches demonstrate that RL policies can benefit from human-like visual attention, as human gaze often reveals the most salient locations in the environment [211]. While incorporating additional modalities is often advantageous, it inevitably introduces extra complexity into the model design.
@@
Robotics has always raised various ethical, societal, and legal concerns. These include risks related to privacy, job displacement, decision-making bias, and the impact on social norms and human relationships.
@@
Most current VLAs focus on household or industrial settings, but a wider range of applications is possible, such as virtual assistants, autonomous vehicles, and agricultural robots. Various embodiments may also call for specialized VLAs, including dexterous hands, drones, quadruped robots, and humanoid robots. One particularly important field is healthcare, encompassing surgical robots [212], care robots [213]. Healthcare demands higher safety and privacy standards and may necessitate novel techniques such as human-in-the-loop (HITL) control and federated learning. Moreover, due to the significant domain gap, specialized vision models might be needed for medical images.
@@ VII Conclusion
Vision-language-action models hold immense promise for enabling embodied agents to interact with the physical world and fulfill users’ instructions. This paper is the first survey to review large VLAs alongside generalized VLAs. Our taxonomy provides a high-level overview of three main lines of research: key components, control policies, and task planners. We meticulously analyze and compare their technical details, including model architectures, training strategies, and individual modules. Additionally, we highlight essential resources for training and evaluating VLAs, such as datasets, simulators, and benchmarks. We hope this survey captures the rapidly evolving landscape of embodied AI and inspires future research.